Fifty grams of Chlorine reacts completely with Phosphorus obtaining ninety grams of the product.


P4(s) + Cl2(g) = PCl3(l) 


a. Balance the chemical equation;

b. Identify the limiting reactant and the excess reactant if applicable;

c. Compute for the theoretical yield; d. Determine the percent yield of the reaction;

e. Calculate the percent error and

f. Compute for the excess amount of the excess reactant if applicable.

g. Show the complete solutions for your answers in a separate paper.


1
Expert's answer
2021-04-14T06:10:07-0400



Balanced :

P4 + 6Cl2 → 4PCl3


Moles of Cl2 "= \frac{50}{35}=1.43moles"


Moles of product"(PCl_3)=\frac{4×1.43}{6}=0.95 moles"



Limiting reactant is "PCl_3" and the excess reactant is "Cl_3"


Theoretical yield "= 0.95\u00d7 137= 130.15 g"


Percent yield "= \frac { 90}{130.15}×100=69.15"%